e^-t/4.6=1.73^2.7t The answer is t=0, how do you get this?
Answer by scott8148(6628)

if t=0, then the exponents are zero and (by definition) both sides equal one

this is a "default" solution ___ changing any or all other values in the equation does not change the solution



___ alternatively ___

taking the natural log ___ -t / 4.6 = ln(1.73) * 2.7t

multiplying by 4.6 ___ -t = ln(1.73) * 12.42t

adding t-ln(1.73) ___ -ln(1.73) = 13.42t

dividing by 13.42 ___ -ln(1.73) / 13.42 = t

t = -.0408 (approx)